OpenStudy (anonymous):

Two years ago, Henry weighed 30 lb more than Walter. Since that time, Walter had lost 10 lb, while Henry's weights stayed the same. Their present weight together is 300 lb, find how many pounds each weighs now.

OpenStudy (anonymous):

2 years ago: H=W+30. Now W is 10 less, so replace W with W-10 to get H=(W-10)+30. We know H=300, so solve for W.

OpenStudy (anonymous):

wait, sorry their weight together is 300

OpenStudy (anonymous):

Oh sorry, I misread the question. So we know W+H=300 and H=(W-10)+30. Can you solve these equations simultaneously?

OpenStudy (anonymous):

uh... if i can simplify the terms? one second

OpenStudy (anonymous):

w+110 and h=140

OpenStudy (anonymous):

Not quite - yours don't add up to 300 :) H=W+20 W+H=W+W+20=300 2W+20=300 W=140 H=W+20 W=140+20=160

OpenStudy (anonymous):

oh lol